HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

OBSERVATION OF THE B^{6}Π STATE OF CrH IN A VERNIER-FILTERED FREQUENCY COMB ABSORPTION EXPERIMENT

J. MORVILLE, \underline{P. CROZET} , \underline{A. J. ROSS}, Université de Lyon, Université Claude

Bernard Lyon 1 & CNRS, Institute Lumière Matière UMR 5309, 69622 Villeurbanne, France

A cavity-enhanced absorption spectrum of CrH, recorded using a mode-locked Ti:sapphire laser as a light source, shows hitherto unassigned (and irregularly spaced) features around 775 nm, in the region of the (1–0) P branches of the A ^{6}Σ^{+}–X ^{6}Σ^{+} system. They are indicated with dots in the figure shown below, and are tentatively assigned as B ^{6}Π_{7/2}–X ^{6}Σ^{+} transitions. These assignments place the 7/2 spin component of the B state about 12900 cm^{−1} above the lowest level of the electronic ground state.

Figure 1: Marked features do not belong to the A-X (1-0) band of CrH, and are assigned as P_{1} lines of the B-X system.

CrH was formed in a DC discharge sputter source, fitted with a pierced cathode made of iron/chromium alloy. A continuous flow of H_{2} in argon passed through the cathode into a vacuum chamber with high reflectance mirrors mounted on its sidearms. Placed at the centre of this high-finesse (F=16000) optical cavity, the source produced a plume about 1 cm wide, which was probed by the laser beam. Transmission through the optical cavity was recorded on a split photodiode, using a Vernier filtering scheme^{1} to achieve fast, broadband coverage close to Doppler resolution. Several spectra were averaged and concatenated to cover the region 12500–13100 cm^{−1}, providing relative intensity information. Some lines were also recorded in cw laser excitation, offering a cleaner baseline and giving some reference calibration for the Vernier spectrum, as part of a search for weaker (satellite) transitions needed to confirm our assignments.

1Lu et al Photonics \underline{9} 222 (2022)

p-number: p087  Submitted on Mon, 12 Jun 2023 16:43:53 +0200